( * ) If x = a(cosec a + cot a) and y = b(1 – cos a) / sin a, then

( * )
If  x = a(cosec a + cot a) and y = b(1 – cos a) / sin a, then

Grade:11

1 Answers

Nishant Vora IIT Patna
9 years ago
Hello student, Please find the answer to your question below

x = a(cosec a + cot a)
x = a(1/sina + cosa / sin a)
x = a(1+cosa / sin a)
and y = b(1 – cos a) / sin a

From here i believe the question is incomple
what parrt is to be proved is not mentioned in the ques